Circular No. 9093

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V496 SCUTI = NOVA SCUTI 2009 S. Nakano, Sumoto, Japan, reports the discovery by H. Nishimura

(Miyawaki, Kakegawa) of a possible nova on two 10-s CCD frames taken on Nov. 8.370 UT using a Canon EOS 5D camera (+ Minolta 120-mm lens); Nakano has measured mag 8.8 and position α = 18^{h}43^{m}45^{s}.65, δ = −7^{o}36′41′′.5 (equinox 2000.0; uncertainty 2′′.0) for the new object from the JPEG discovery image (limiting mag 12.0). Nothing is visible at this position on frames taken on Nov. 7.377 using the same instrumentation (same limiting mag). E. Guido and G. Sostero obtained unfiltered CCD images (remotely, 0.25-m reflector, GRAS Observatory) on Nov. 9.08, finding the presumed nova to be at mag ∼ 8.5 with position end figures 45^{s}.57, 42′′.0; the extreme stellar crowding due to nearby field stars makes measurement rather difficult, but a Digitized Sky Survey plate (limiting red mag ∼ 20) from 1996 Aug. 13 fails to show a clear pre-outburst precursor; additional details are given on CBET 2008. E. V. Kazarovets reports that the GCVS team has assigned the designation V496 Sct to this object.

D. Balam, Dominion Astrophysical Observatory, National Research Council of Canada (NRCC); and G. Sarty, University of Saskatchewan, write that a spectrogram (range 429–455 nm, resolution 0.02 nm), obtained with the NRCC 1.82-m Plaskett telescope on Nov. 10.08 UT, confirms that this object is a nova, showing Hα emission (HWZI ∼ 400 km/s) with a P-Cyg profile, its absorption minimum being blue-shifted by 700 km/s with respect to the emission peak. Additional lines showing P-Cyg profiles include N II (448.1 nm), Fe II (multiplet 27) at 435.0 and 441.8 nm, and Fe II (multiplet 38) at 452.1 nm.

U. Munari and A. Siviero, Istituto Nazionale di Astrofisica, Padova Astronomical Observatory; and L. Buzzi and P. Valisa, Asiago Novae and Symbiotic Stars (ANS) collaboration, report that low-resolution, absolute spectrophotometry (range 390–860 nm, dispersion 0.21 nm/pixel) of the nova were obtained on Nov. 9.779 UT with the 0.6-m telescope of the Schiaparelli Observatory in Varese. The spectrum is dominated by emission lines due to Balmer, Fe II (multiplets 27, 28, 37, 38, 42, 48, 49, 55, and 74), and O I 777.2- and 844.6-nm, all displaying deep P-Cyg profiles, with the absorption component blue-shifted by 700 km/s from the emission component. The mean FWHM of the emission and absorption of the P-Cyg profiles are 950 and 850 km/s, respectively. The Hα:Hβ:Hγ:Hδ integrated emission-flux ratio is 3.68:1.00:0.23:0.16, and that of O I (777.2 nm)/(844.6 nm) is 0.91. The overall aspect is that of a typical nova of the Fe II class, observed close to maximum brightness.

2009 November 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green